The probability of pulling out a coloured counter from a bag is shown below: Green=0.2. Purple=0.15. Black=0.3. Pink=?. What is the probability of pulling out a pink counter?

When we have a probability question with multiple possible outcome as we have here (the multiple possible outcomes are pulling out a green, purple, black or pink counter) the probabilities of all of these outcome ALWAYS adds up to one. This makes the task of finding the probability of pulling out a pink counter a lot easier! So step one is to add up all of the probabilities that we have been given: 0.2 + 0.15 + 0.3= 0.65Green + Purple + BlackTo work out what the probability of pulling a pink counter would be, all we need to do is 1 minus the total of the three other outcomes:1-0.65=0.35.Therefore, 0.35 is the probability of pulling out a pink counter, if you wanted to double check your answer you could again add up all of the probabilities (including the pink one that you worked out) and it should equal 1!

FA

Related Maths GCSE answers

All answers ▸

What is the gradient of the line passing through the point (1,2) and (5,5)? What is the equation of this line? What is the equation of the line perpendicular to this line that passes through the origin (0,0)?


Given that 346 × 27 = 9342 , work out 34.6 × 2.7


On a packet of brown rice it says 'When 60g of rice is cooked, it will weigh 145g.' If Katy has 100g of brown rice, how much will it weigh when cooked?'


A sequence increases by 5 each time and the first term is x. The sum of the first four terms is 54. Set up and solve an equation to work out the value of x.